OpenStudy (anonymous):

another one lol ummm bet you cant get this one 4/9=x/4 solve each proportion.

OpenStudy (anonymous):

again cross multiply 4*4 = x*9 16/9 = x

OpenStudy (anonymous):

ok i know how to do it now

OpenStudy (anonymous):

good!! do some on your own and tell us the answer to say whether it is correct or not.....